HUBBLE SPACE TELESCOPE DAILY REPORT #1063 PERIOD COVERED: 0700 EST Daily Status Report as of 028/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL/COSTAR 5605 (COSTAR/FOS Fine Alignment X-Tilt/Y-Tilt Repeat of 4747 With New Target) Both the Faint Object Spectrograph and the COSTAR were used to make observations of standard star WD0501+527 to perform alignment and focus operations on the FOS M1 mirror. Observations and stored command moves (X and Y tilt) were monitored when the telemetry was available. All of the guide star acquisitions and reacquisitions were executed as planned. Fine lock was obtained and the peakup operations worked as planned. One observation was reported to have had a recentering event. No other anomalies were noted. 1.2 Completed S/C 5254 (GHRS SMOV Side 1 Activation) The GHRS Side One had low voltage and the DEBs on at 027/1100Z. No problems were noted. The OSS was asked to monitor the MEB temperature (TEMP1). The MEB TEMP1 reached 35.5 degrees before the DEBs were turned off. 1.3 Completed WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Six images were received and displayed. They were reported to be okay. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 1 Successful : 1 Scheduled Reacquisitions : 14 Successful : 14 2.2 FHST Updates: Scheduled :0 Successful :0 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities remain on schedule. GHRS Side 1 activation and low voltage turn-on functions are currently in progress with no reported problems. COSTAR/FOS fine alignment activities continue. There have been no difficulties with any HST subsystem. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 5196 (GHRS SMOV Side 1 High Voltage Turn On) 3.2 Complete WF/PC-2 5549 (GHRS Side 2 Dark Noise Monitor - Cycle 4) 3.3 Complete WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) 3.4 Complete Two Sets of WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) 3.5 Complete Two Sets of WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) 3.6 Complete Two Sets of S/C 5254 (GHRS SMOV Side One Activation) 3.7 Complete WF/PC-2 4749 (WF/PC-2 Relative Photometric Calibration) 3.8 Complete COSTAR/U 5605 (COSTAR/FOS Fine Alignment X-Tilt/Y-Tilt (Repeat of 4747 With A New Target)) 3.9 Complete FOC/96 5152 (Shape and Expansion of SN1987A Ejecta) 3.10 Complete WF/PC-2 5457 (A Snapshot Survey for Blue Straggles in Magellanic Cluster Cores) 3.11 Complete S/C 4435 (Spacecraft Disturbance Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-